Of 250 adults who tried a new product,187 rated it excellent. of 100 children sampeld, 66 rated it excellent. test the hypothesis that the two population proportions are equal using 0.01 level.
Answer by stanbon(75887)

Of 250 adults who tried a new product,187 rated it excellent. of 100 children sampeld, 66 rated it excellent. test the hypothesis that the two population proportions are equal using 0.01 level.
--------------------------
p1-hat = 187/250 = 0.75
--
p2-hat = 66/100 = 0.67
---------------------------
Ho: p1-p2 = 0 (claim)
Ha: pa-p2 # 0
-------
sample proportion: p1-p2 = 0.75-0.67 = 0.08
-----
test stat: z(0.08) = 0.08/sqrt[(0.75*0.25/250)+(0.67*0.33/100)] = 1.4702
----
p-value = 2*P(z >1.4702) = 0.1416
----
conclusion: Since the p-value is greater than 1%, fail
to reject Ho.
The test results support the claim.
===================
Cheers,
Stan H.